OpenStudy (anonymous):

Find what amount 125 is 25% of using the decimal approach.

Parth (parthkohli):

Hmm I see. Equate here. \(\Large \color{Black}{\Rightarrow 125 = 0.25x}\) \(\Large \color{Black}{\Rightarrow {125 \over 0.25} = x }\) \(\Large \color{Black}{\Rightarrow 125 \times 4 =x }\) \(\Large \color{Black}{\Rightarrow x = 500 }\)

Parth (parthkohli):

25% is the same thing as 0.25

Parth (parthkohli):

If you ask why, then see, 25% = 25/100 = 0.25

OpenStudy (anonymous):

Wow! Thank you loads!! :D I really need to know this.